OpenStudy (anonymous):

Find the surface area of the cylinder in terms of pi. A) 396 pi in^2 B) 522 pi in^2 C) 360 pi in^2 D) 1008 pi in^2

OpenStudy (mathstudent55):

Shouldn't we be given some dimensions?

OpenStudy (anonymous):

sorry forgot the picture

OpenStudy (mathstudent55):

For the bases, upper and lower surface, you have circles with radius = 9 For the side, you have a rectangle in which one side is the height, 20 inches, and the other side is the circumference of the base. Abase = (pi)r^2 Aside = 2(pi)r Remember that these are 2 bases, so Total Area = 2(Abase) + Aside = 2(pi)r^2 + 2(pi)rh = 2(pi)(9^2) + 2(pi)(9)(20) = 162(pi) + 360(pi) = 522(pi) in.^2

OpenStudy (anonymous):

woohoo! you rock thx so much

OpenStudy (mathstudent55):

The bases are circles, so the areas are easy to see. Think of the lateral (side) area as if you peeled off the label of a can and laid it flat. The shape is a rectangle. The width of that rectangle is the height of the can. The length of the rectangle is wrapped around the can, so it's the circumference of the circle which is the base.

OpenStudy (anonymous):

okay thx (: